Find the next two
numbers in the
sequence. What is
the rule?

100, 25, 50, 12.5, 25,
____, ____

Answer by richwmiller(17219)

One series could be
A(n+1)=An*1/4 then A(n+2)=A(n+1)*2
100*1/4=25
25*2=50
50*1/4=12.5
12.5*2=25
25*1/4=6.25
6.25*2=12.5
so the next two could be
6.25 and 12.5
Another possibility
100 -75 100 -162.5 275
25 25 -62.5 112.5 275
50 -37.5 50 387.5
12.5 12.5 437.5
25 450.0
475
The sixth term could be 475
100 -75 100 -162.5 275 0
25 25 -62.5 112.5 275 0
50 -37.5 50 387.5 275
12.5 12.5 437.5 662.5
25 450 1100
475 1550
2025
7th term could be 2025
so the 6th and 7th terms could be 475 and 2025